OpenStudy (anonymous):

What is the slope of the line that passes through the points (–1, 13) and (4, –2)?

OpenStudy (anonymous):

i got this...

OpenStudy (anonymous):

hehe -3

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

np :) here to help

OpenStudy (the_fizicx99):

The slope formula is M = y2-y1/x2-x1 -2 - 13 / 4 - (-1) = -15/5 = -3 M = -3

OpenStudy (anonymous):

thanks Ill definitely use that next time and record it @tHe_FiZiCx99

OpenStudy (the_fizicx99):

Yw, you can't give away direct answers though, it violates the CoC >,<